-3 -4 If A= Then A-1 = 5 6 Select one: 3 a. -5 2 2 2 W 3 2 5 2 O b. 2. 3 2 O c. 3 5 2 3 d. 4 6
Answer #1

Ama a Note for ad=bc to el с g Ja aloc and to the o th A -3 G 6 ts th - ad-bc=-3x6 (-4x52 -18 + 20 2 thu 6. - 1 A 4 = 2 -5 -3